# Efficient Diffusion Model for Image Restoration by Residual Shifting

### v1

@@ III-C Discussion on Arbitrary Resolution
To analyze the underlying reason, we visualize the attention maps extracted from the first attention layer in this baseline network, as shown in Fig. 4(c1) and (c2). Note that these two attention maps are both interpolated to the resolution of $256\times 256$ for ease of comparison. Evidently, the example with a larger resolution tends to generate a more uniformly distributed attention map, i.e., Fig. 4(c2), being consistent with the entropy values annotated on the left-upper corner [^2] The principle of maximum entropy posits that it achieves the maximum entropy when the attention map conforms to a uniform distribution. Consequently, a uniformly distributed attention map often leads to an over-smooth outcome, introducing undesirable distortions in performance.
To address this issue, some recent studies [14, 12] have chosen to discard the self-attention layers, a strategy that typically results in a noticeable decline in performance. Inspired by Liang et al. [40], we propose a solution by substituting the self-attention layer with that in Swin Transformer [80]. This straightforward replacement not only alleviates the blurring problem but also maintains the promised performance, as shown in Fig. 4(d1) and (d2). This is because the Swin Transformer computes the attention map in a local window, thus being independent of the image resolution.
@@ IV Experiments on Image Super-resolution
This section offers an evaluation of the proposed method on the task of image super-resolution (SR), with a particular focus on the setting of ×$\times 4$ SR following existing studies [39, 38]. We first provide an in-depth analysis on the proposed model and then conduct a thorough comparison against recent state-of-the-art methods (SotAs) on one synthetic and two real-world datasets. For the sake of brevity in presentation, our method is herein referred to ResShift or ResShiftL. The former is trained based on the primary loss of diffusion model in Eq. (10), while the latter further introduces the perceptual regularization as shown in Eq. (11). To facilitate a more profound understanding of our designed diffusion model and the noise schedule, the model analysis part of Sec. IV-B mainly centers around ResShift.
Figure 5: Qualitative comparisons of ResShift under different combinations of ($T$, $p$, $\kappa$) on the task of image super-resolution. For example, “(15, 0.3, 2.0)” represents the recovered result with $T=15$, $p=0.3$, and $\kappa=2.0$. Please zoom in for a better view.
TABLE I: Performance comparison of ResShift on the ImageNet-Test dataset for image super-resolution under different configurations.
  Configurations Metrics
$T$
$p$
$\kappa$
[…]
4
0.3 2.0
25.64
0.6903
0.3242
10
25.20
0.6828
0.2517
15
25.01
0.6769
0.2312
30
24.52
0.6585
0.2253
50
24.22
0.6483
0.2212
15
0.3
2.0
25.01
0.6769
0.2312
0.5
25.05
0.6745
0.2387
1.0
25.12
0.6780
0.2613
2.0
25.32
0.6827
0.3050
3.0
25.39
0.5813
0.3432
15 0.3
0.5
24.90
0.6709
0.2437
1.0
24.84
0.6699
0.2354
2.0
25.01
0.6769
0.2312
8.0
25.31
0.6858
0.2592
16.0
24.46
0.6891
0.2772
@@ IV-A Experimental Setup
Figure 6: Perception-distortion trade-off of ResShift and LDM. The vertical and horizontal axes represent the strength of the perception and distortion, measured by LPIPS and MSE, respectively.
@@
where $\bm{k}$ is the blurring kernel, $\bm{n}$ is the noise, $\bm{y}$ and $\bm{x}$ denote the LQ image and HQ image, respectively. To comprehensively evaluate the performance of our model, we consider more complicated types of blurring kernel, downsampling operator, and noise type. The detailed settings on them can be found in Appendix -C1. It should be noted that we select the HQ images from ImageNet [81] instead of the prevailing datasets in SR, such as Set5 [85], Set14 [86], and Urban100 [87]. The rationale behind this setting is rooted in the fact that these prevailing datasets only contain very few source images, which fail to thoroughly evaluate the performance of various methods under a multitude of different degradation types.
Two real-world datasets are adopted to evaluate the efficacy of our method. The first is RealSR-V3 [88], containing 100 real images captured by Canon 5D3 and Nikon D810 cameras. Additionally, we collect another real-world dataset named RealSet80. It comprises 50 LR images widely used in recent literature [89, 90, 91, 92, 38, 93]. The remaining 30 images are downloaded from the internet by ourselves.
@@
Methods
Metrics
@@ Methods
Runtime (s)
#Params (M)
ESRGAN [72]
20.67
0.448
0.485
0.038
16.70
RealSR-JPEG [94]
23.11
0.591
0.326
0.038
16.70
BSRGAN [39]
24.42
0.659
0.259
0.038
16.70
SwinIR [40]
23.99
0.667
0.238
0.107
28.01
RealESRGAN [38]
24.04
0.665
0.254
0.038
16.70
DASR [95]
24.75
0.675
0.250
0.022
8.06
LDM-15 [25]
24.89
0.670
0.269
0.247
[…]
LDM-4 [25]
24.74
0.657
0.345
0.077
StableSR-15 [29]
23.37
0.631
0.262
1.070
[…]
StableSR-4 [29]
24.11
0.658
0.287
0.399
ResShift-15
25.01 0.677 0.231
0.682
[…]
ResShiftL-4
25.02 0.683 0.208
0.186
Figure 7: Qualitative results of different methods on the synthetic ImageNet-Test dataset for image super-resolution. Note that we only display the comparison results to the recent five SotA methods in (b)-(f) due to the page limitation, and the whole results can be found in Fig. 14 of the appendix. Please zoom in for a better view.
TABLE III: Quantitative results of different methods on two real-world datasets for image super-resolution. Note that the results of diffusion-based methods are denoted as “Method-A”, where “A” represents the number of sampling steps. The best and second best results are highlighted in bold and underline.
Methods
Datasets
@@
ESRGAN [72]
0.2362
29.048
0.4165
48.153
RealSR-JPEG [94]
0.3615
36.076
0.5828
57.379
BSRGAN [39]
0.5439
63.586
0.6263
66.629
SwinIR [40]
0.4654
59.632
0.6072
64.739
RealESRGAN [38]
0.4898
59.678
0.6189
64.496
DASR [95]
0.3629
45.825
0.5311
58.974
LDM-50 [25]
0.4907
54.391
0.5511
55.826
LDM-15 [25]
0.3836
49.317
0.4592
50.972
StableSR-50 [29]
0.5208
60.177
0.6214
62.761
StableSR-15 [29]
0.4974
59.099
0.5975
61.476
ResShift-15
0.6014
58.979
0.6645
62.782
ResShiftL-4 0.5854
56.184
0.6483
61.017
Compared Methods. We evaluate the effectiveness of ResShift and ResShiftL in comparison to seven recent SR methods, namely RealSR-JPEG [94], BSRGAN [39], RealESRGAN [38], SwinIR [40], DASR [95], LDM [25], and StableSR [29]. LDM and StableSR are both diffusion-based methods trained with 1,000 diffusion steps. For a fair comparison, we accelerate them to 15 or 4 steps using the DDIM [36] algorithm during inference in this study. The hyper-parameter $\eta$ in DDIM is set to be 1.0 as this value yields the most realistic results. For clarity, the results of LDM and StableSR are denoted as “Method-A”, where “A” represents the number of inference steps.
Evaluation Metrics. We evaluate the efficacy of different methods using three widely used metrics in most of the SR literature, namely PSNR, SSIM [96], and LPIPS [77], on the synthetic dataset. Notably, PSNR and SSIM are calculated in the luminance (Y) channel of YCbCr space, while LPIPS is directly computed in the standard RGB (sRGB) space. For assessments on real-world datasets, we mainly rely on CLIPIQA [97] and MUSIQ [98] as evaluative metrics to compare the performance of various methods. CLIPIQA and MUSIQ are both non-reference metrics specifically designed for assessing the realism of images. Of particular note, CLIPIQA leverages the CLIP [99] model, pre-trained on the extensive Laion400M [100] dataset, thereby demonstrating strong generalization ability.

### v2

@@ III-D Discussion on Arbitrary Resolution
To analyze the underlying reason, we visualize the attention maps extracted from the first attention layer in this baseline network, as shown in Fig. 4(c1) and (c2). Note that these two attention maps are both interpolated to the resolution of $256\times 256$ for ease of comparison. Evidently, the example with a larger resolution tends to generate a more uniformly distributed attention map, i.e., Fig. 4(c2), being consistent with the entropy [^2] The average entropy of the attention map $\bm{W}\in\mathcal{R}^{n\times n}$ is defined as $-\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{n}w_{ij}\ln{w_{ij}}$, where we assume that each row of $\bm{W}$ represents the event probabilities of a discrete categorical distribution. values annotated on the left-upper corner [^3] The principle of maximum entropy posits that it achieves the maximum entropy when the attention map conforms to a uniform distribution. Consequently, a uniformly distributed attention map often leads to an over-smooth outcome, introducing undesirable distortions in performance.
To address this issue, some recent studies [4, 14] have chosen to discard the self-attention layers, a strategy that typically results in a noticeable decline in performance. Inspired by Liang et al. [41], we propose a solution by substituting the self-attention layers with Swin Transformers [44]. This straightforward replacement not only alleviates the blurring problem but also maintains the promised performance, as shown in Fig. 4(d1) and (d2). This is because the Swin Transformer computes the attention map in a local window, thus being independent of the image resolution.
@@ IV Experiments on Image Super-resolution
This section offers an evaluation of the proposed method on the task of image super-resolution (SR), with a particular focus on the setting of ×$\times 4$ SR following existing studies [40, 39]. We first provide ablation studies of the proposed model and then conduct a thorough comparison against recent state-of-the-art methods (SotAs) on one synthetic and two real-world datasets. For brevity in presentation, our method is herein referred to ResShift or ResShiftL. The former is trained based on the primary loss in Eq. (10) with 15 diffusion steps, while the latter further introduces the perceptual regularization as shown in Eq. (11) with 4 steps.
@@ IV-A Experimental Setup
where $\bm{k}$ is the blurring kernel, $\bm{n}$ is the noise, $\bm{y}$ and $\bm{x}$ denote the LQ image and HQ image, respectively. To comprehensively evaluate the performance of our model, we consider more complicated types of blurring kernel, downsampling operator, and noise type. The detailed settings on them can be found in Appendix -B1. It should be noted that we select the HQ images from ImageNet [91] instead of the prevailing datasets in SR, such as Set5 [95], Set14 [96], and Urban100 [97]. The rationale behind this setting is that these datasets only contain very few source images, which fail to thoroughly evaluate the performance of various methods under many different degradation types.
TABLE I: Quantitative comparisons of the proposed method with different attention layers on the synthetic dataset of ImageNet-Test and the real-world dataset of RealSet80.
  Methods Attention types ImageNet-Test RealSet80
9
@@
  Baseline Self-attention 24.97 0.6806 0.2137 0.5934 51.844 0.5883 59.090
ResShiftL Swin Transformer 25.02 0.6833 0.2076 0.5976 51.966 0.6418 61.022
TABLE II: Quantitative comparisons of our method with various fidelity losses ($L_{1}$ or $L_{2}$) on the ImageNet-Test dataset.
  Methods Metrics
@@
  ResShiftL-L1 24.63 0.6710 0.2115 0.6261 53.8254
ResShiftL 25.02 0.6833 0.2076 0.5976 51.9656
TABLE III: Quantitative comparisons of our method with ($\lambda=1$) or without ($\lambda=0$) perceptual loss on the ImageNet-Test dataset.
  Hyper-parameters Metrics
6
@@
  $\lambda=0$ 25.64 0.6930 0.3242 0.4241 41.8308
$\lambda=1$ 25.02 0.6833 0.2076 0.5976 51.9656
Two real-world datasets are adopted to evaluate the efficacy of our method. The first is RealSR-V3 [98], containing 100 real images captured by Canon 5D3 and Nikon D810 cameras. Additionally, we collect another real-world dataset named RealSet80. It comprises 50 LQ images widely used in recent literature [99, 100, 101, 102, 39, 81]. The remaining 30 images are downloaded from the internet by ourselves.
Figure 5: Ablation studies of our method regarding the perceptual loss.
TABLE IV: Quantitative results and the corresponding standard deviation (std) of the proposed method under multiple random seeds on the dataset of ImageNet-Test.
  Metrics Seed-1 Seed-2 Seed-3 Seed-4 Std
  PSNR $\uparrow$ 25.02 25.01 25.03 25.01 0.00829
SSIM $\uparrow$ 0.6833 0.6826 0.6834 0.6830 0.00031
LPIPS $\downarrow$ 0.2076 0.2074 0.2076 0.2075 0.00008
TABLE V: Quantitative comparison on performance, running time, and the number of parameters of different methods on ImageNet-Test dataset for Image Super-resolution. The results of the diffusion-based methods are denoted as “Method-A”, where “A” represents the number of sampling steps. Running time is tested on NVIDIA Tesla A100 GPU on the x4 (64$\rightarrow$256) SR task. The non-trainable parameters, such as the parameters of VQGAN in LDM, are marked with gray color for clarity.
  Methods Metrics
8
@@
  ESRGAN [75] 20.67 0.448 0.485 0.451 43.615 0.038 16.70
RealSR-JPEG [103] 23.11 0.591 0.326 0.537 46.981 0.038 16.70
BSRGAN [40] 24.42 0.659 0.259 0.581 54.697 0.038 16.70
SwinIR [41] 23.99 0.667 0.238 0.564 53.790 0.107 28.01
RealESRGAN [39] 24.04 0.665 0.254 0.523 52.538 0.038 16.70
DASR [104] 24.75 0.675 0.250 0.536 48.337 0.022 8.06
DiffIR-4 [27] 24.50 0.674 0.217 0.554 54.567 0.161 26.48
  LDM-50 [25] 24.17 0.637 0.245 0.600 52.665 0.773 113.60+55.32
LDM-15 [25] 24.89 0.670 0.269 0.512 46.419 0.247
LDM-4 [25] 24.74 0.657 0.345 0.372 38.161 0.077
  StableSR-50 [30] 22.96 0.611 0.264 0.666 59.559 3.205 52.49+1422.49
StableSR-15 [30] 23.37 0.631 0.262 0.660 59.492 1.070
StableSR-4 [30] 24.11 0.658 0.287 0.580 53.698 0.399
  CCSR-45 [42] 24.67 0.661 0.236 0.614 58.242 4.500 363.15+1303.60
CCSR-15 [42] 24.86 0.669 0.243 0.581 55.773 1.670
CCSR-4 [42] 25.37 0.694 0.282 0.450 46.204 0.622
  ResShift-15 25.01 0.677 0.231 0.592 53.660 0.682 118.59+55.32
ResShiftL-4 25.02 0.683 0.208 0.598 51.966 0.186
Figure 6: Qualitative results of different methods on the synthetic ImageNet-Test dataset for image super-resolution. Note that we only display the comparison results to the recent five SotA methods in (b)-(f) due to the page limitation, and the complete results are presented in Fig. 18 of the Appendix.
TABLE VI: Quantitative results of different methods on two real-world datasets for image super-resolution. Note that the results of diffusion-based methods are denoted as “Method-A”, where “A” represents the number of sampling steps.
  Methods Datasets
@@
  ESRGAN [75] 0.2362 29.048 0.4165 48.153
RealSR-JPEG [103] 0.3615 36.076 0.5828 57.379
BSRGAN [40] 0.5439 63.586 0.6263 66.629
SwinIR [41] 0.4654 59.632 0.6072 64.739
RealESRGAN [39] 0.4898 59.678 0.6189 64.496
DASR [104] 0.3629 45.825 0.5311 58.974
DiffIR-4 [27] 0.4315 57.449 0.5909 62.028
  LDM-50 [25] 0.4907 54.391 0.5511 55.826
LDM-15 [25] 0.3836 49.317 0.4592 50.972
LDM-4 [25] 0.2865 43.205 0.3582 45.182
  StableSR-50 [30] 0.5208 60.177 0.6214 62.761
StableSR-15 [30] 0.4974 59.099 0.5975 61.476
StableSR-4 [30] 0.4392 56.179 0.5250 57.445
  CCSR-45 [42] 0.5681 63.222 0.6385 65.889
CCSR-15 [42] 0.5540 62.331 0.6284 64.859
CCSR-4 [42] 0.4893 58.039 0.5550 59.646
  ResShift-15 0.5958 58.475 0.6645 62.782
ResShiftL-4 0.5995 57.554 0.6418 61.022
Compared Methods. We evaluate the effectiveness of ResShift and ResShiftL in comparison to nine recent SR methods, namely RealSR-JPEG [103], BSRGAN [40], RealESRGAN [39], SwinIR [41], DASR [104], LDM [25], DiffIR [27], StableSR [30], and CCSR [42]. For a fair comparison, we accelerate the diffusion-based methods, including LDM, DiffIR, StableSR, and CCSR, to 15 or 4 steps using their default accelerating algorithm during inference. For clarity, the results of these diffusion-based methods are denoted as “Method-A”, where “A” represents the number of inference steps.
Evaluation Metrics. We evaluate the efficacy of different methods using five widely used metrics, including PSNR, SSIM [105], LPIPS [84], CLIPIQA [106], and MUSIQ [107]. Notably, CLIPIQA and MUSIQ are both non-reference metrics specifically designed for assessing the realism of images. Particularly, CLIPIQA leverages the CLIP [108] model, pre-trained on the extensive Laion400M [109] dataset, thereby demonstrating strong generalization ability.
Figure 7: Qualitative comparisons on three real-world examples from RealSet80. Please zoom in for a better view.
IV-B Ablation Studies
In this part, we provide some necessary ablation studies on several components in our model. More comprehensive analysis about the noise schedule, perception-distortion trade-off, and comparisons with more advanced samplers can be found in Appendix -B2.
Fidelity loss. The loss function of our method incorporates both a fidelity loss and a perceptual regularizer, as shown in Eq. (11). The fidelity loss is formulated as the $L_{2}$ norm, quantifying the discrepancy between the predicted HQ image and the underlying ground truth. We have also explored the use of an $L_{1}$ norm in place of the $L_{2}$ norm for the fidelity loss, resulting in a variant of our model denoted as ResShiftL-L1. Comparison results are summarized in Table II, demonstrating that ResShiftL outperforms ResShiftL-L1 on reference metrics, while ResShiftL-L1 shows superior performance on non-reference metrics. Considering the high fidelity requirement for IR, we adopted the $L_{2}$ norm in this study.
Perceptual loss. In contrast to our conference version [43], this study integrates an additional perceptual regularizer, detailed in Eq. (11), which enhances the model efficiency by reducing the sampling steps from 15 to 4. The ablation study summarized in Table III indicates that while the introduction of the perceptual loss results in a slight decrease in PSNR and SSIM, it yields significant improvements in LPIPS, CLIPIQA, and MUSIQ. These latter three metrics more truthfully reflect the perceptual quality and realism of images, as supported by the visual comparisons in Fig. 5. Therefore, considering both performance and efficiency, the incorporation of the perceptual regularizer proves to be a critical enhancement.
Swin Transformer. As discussed in Sec. III-D, we replace the self-attention layers in the diffusion Unet with Swin Transformer blocks to address the arbitrary resolution issue. Table I provides a quantitative comparison of this modification. On the synthetic ImageNet-Test dataset, where both training and testing images are of consistent resolution, models with either self-attention layers or Swin Transformer blocks exhibit comparable performance. In contrast, on the real-world dataset RealSet80, which contains images of varying resolutions, the baseline model using self-attention layers suffers from a significant performance drop. This is mainly attributed to the inability of self-attention layers to generalize across resolutions that largely deviate from those encountered during training. A more comprehensive analysis and visualization from the perspective of information entropy are presented in Sec. III-D and Fig. 4.
Sampling Randomness. We discuss the sampling randomness caused by the stochastic sampling of diffusion models within the task of IR. Firstly, IR is an ill-posed problem, particularly in severely degraded scenarios where multiple HQ outputs can correspond to a single LQ image. The random sampling mechanism of diffusion models facilitates a one-to-many mapping, effectively addressing this ill-posed issue by generating diverse but plausible restoration outcomes for any testing image. Secondly, high fidelity is crucial for the task of IR. Our proposed method designs a diffusion process between the HQ and LQ images, rather than starting from random Gaussian noise, which reduces the randomness of sampling to a certain extent. Additionally, Table IV lists the quantitative comparisons of our method under various random seeds, and corresponding visual results can be found in Fig. 20 of the Appendix. These results empirically demonstrate the consistency across different outputs. On the other hand, while some level of randomness is present, it is manageable and beneficial for handling the ill-posedness of IR.
TABLE VII: Quantitative comparisons of various methods on the test dataset ImageNet-Test for inpainting. The best and second best results are highlighted in bold and underline, respectively.
  Mask Types Metrics Methods
9
DeepFillv2 [5] LaMa [6] RePaint [18] DDRM [28] Score-SDE [21] MCG [29] ResShiftL
@@ IV-B Ablation Studies
[…]
TABLE VIII: Quantitative comparisons of various methods on the test dataset CelebA-Test for inpainting. The best and second best results are highlighted in bold and underline, respectively.
  Mask Types Metrics Methods
9
DeepFillv2 [5] LaMa [6] RePaint [18] DDRM [28] Score-SDE [21] MCG [29] ResShiftL
[…]
IV-C Evaluation on Synthetic Data
We present a comparative analysis of the proposed method with recent SotA approaches on the ImageNet-Test dataset, as summarized in Table V and Fig. 6. This evaluation reveals the following conclusions: i) Diffusion-based methods demonstrate significant advantages in terms of non-reference metrics; however, their performance on reference metrics is hindered by the inherent randomness in the sampling procedure. ii) Among diffusion-based methods, our proposed method exhibits superior performance across both reference and non-reference metrics with the same number of sampling steps, indicating an improved fidelity-realism trade-off. iii) ResShiftL is notably faster than other diffusion-based methods, achieving a preeminent balance between performance and efficiency. Even in comparison with the SotA GAN-based method SwinIR [41], it not only maintains comparable speed but also delivers superior performance. This efficiency is attributed to our well-designed diffusion model, which has a shorter transition trajectory.
IV-D Evaluation on Real-World Data
Table VI lists the comparative evaluation using CLIPIQA [106] and MUSIQ [107] for various approaches on two real-world datasets, namely RealSR-V3 [98] and RealSet80. Note that CLIPIQA, benefiting from the powerful representative capability inherited from CLIP, performs consistently and robustly in assessing the perceptional quality of natural images. The results in Table VI reveal that the proposed ResShift or ResShiftL notably outperforms existing methods in terms of CLIPIQA. This suggests that the restored outputs by our method better align with human visual and perceptive systems. In the case of MUSIQ evaluation, ResShift attains competitive performance when compared to current SotA methods, namely BSRGAN [40], SwinIR [41], and RealESRGAN [39]. Collectively, our method shows promising capability in addressing real-world SR challenges.
We display three real-world examples in Fig. 7. We consider diverse scenarios, including text, animal, and natural images to ensure a comprehensive evaluation. An obvious observation is that ResShift or ResShiftL produces more naturalistic image structures. We note that the recovered results of LDM [25] and StableSR [30] are excessively smooth when compressing the inference steps to match with our proposed method, i.e., 15 or 4 steps, largely deviating from the training procedure’s 1,000 steps. Even though other GAN-based methods may also succeed in hallucinating plausible structures to some extent, they are often accompanied by obvious artifacts.
Figure 8: Visual comparisons of various methods on the test dataset of ImageNet-Test for inpainting. The results of diffusion-based methods are denoted as “Method-A”, where “A” represents the number of sampling steps. The masked areas are highlighted using a purple color. Please zoom in for a better view.
V Experiments on Image Inpainting
The proposed diffusion model is a general framework for IR. This section presents a series of experiments to validate its effectiveness in the task of image inpainting. Additional experimental results on blind face restoration and image deblurring are provided in Appendix -C.
V-A Experimental Setup
Training Details. In addressing the task of inpainting, we train two variants of the ResShiftL model, both implemented at a resolution of $256\times 256$. These two variants are tailored for natural images and facial images, respectively. The former model is trained using the training dataset of ImageNet [91], while the latter is trained on the widely used face dataset FFHQ [110]. During training, we randomly generate the image masks to synthesize the LQ images following LaMa [6]. The other training configurations are kept consistent with those in image super-resolution, as detailed in Sec. IV-A.
Figure 9: Visual comparisons of various methods on the test dataset of CelebA-Test for inpainting. The results of diffusion-based methods are denoted as “Method-A”, where “A” represents the number of sampling steps. The masked areas are highlighted using a purple color. Please zoom in for a better view.
Figure 10: One typically failed example on the task of natural image inpainting.
Test Datasets. Two test datasets are constructed by randomly selecting 2,000 images from the validation dataset of ImageNet [91] and CelebA-HQ [111], to facilitate an assessment on natural images and facial images, respectively. These images in each dataset are uniformly divided into four groups to synthesize different types of masked images. To ensure a thorough evaluation, four distinct mask types, denoted as “Box” mask, “Irregular” mask, “Half” mask, and “Expand” mask, are considered as visually shown in Fig. 8. For each mask type, we randomly generate a set of 500 masks, and then employ them to simulate the LQ images. These two datasets are denoted as ImageNet-Test and CelebA-Test in this section.
Compared Methods. In order to evaluate the efficacy of ResShiftL, a comparative analysis is conducted against two GAN-based methods, including DeepFillv2 [5] and LaMa [6], as well as four diffusion-based methods, namely Score-SDE [21], RePaint [18], DDRM [28], and MCG [29]. For the diffusion-based methods, we accelerate their sampling process to 250 steps using the DDIM [37] algorithm.
Evaluation Metrics. For the sake of comprehensively assessing the performance of various approaches, we adopt one full-reference metric LPIPS [84] and one no-reference metric CLIPIQA [106] as our principal evaluative criteria.
V-B Comparison with SotA Methods
We provide a quantitative evaluation of different methods on the test dataset of ImageNet-Test and CelebA-Test, as detailed in Table VII and Table VIII, respectively. The proposed ResShiftL achieves the best or, at the very least, comparable performance to recent SotA methods across most cases, particularly excelling in the more challenging mask types such as “Irregular”, “Half”, and “Expand”. In comparison to other diffusion-based approaches, ResShiftL still maintains a competitive advantage, even with a significantly reduced number of sampling steps (4 vs. 250).
A series of visual illustrations on various mask types are displayed in Fig. 8 and Fig. 9. In the case of “Box” mask, recent methods, namely LaMa [6] and MCG [29], and our ResShiftL all perform well. For the other three mask types containing large occluded areas, most of the comparison methods fail to handle such complicated scenarios. In contrast, the proposed ResShiftL consistently yields more plausible and realistic results under these scenarios, especially on the preservation of coherency to the unmasked regions. The qualitative analysis presented herein reaffirms the stability and exceptional performance of ResShiftL, aligning with the quantitative comparison above.
While ResShiftL has demonstrated strong performance in most scenarios, failed examples still exist, particularly in cases involving large masked areas, as illustrated in Fig. 10. Existing methods struggle to effectively deal with such an extremely occluded example, mainly because the available information in this image is too limited. To address this challenge, a potential improvement avenue is introducing more supplementary guidance, such as text prompts. We leave the exploration in this direction for future research.
VI Conclusion
In this work, we have introduced an efficient diffusion model specifically designed for IR. Unlike existing diffusion-based IR methods that require a large number of iterations to achieve satisfactory results, our proposed method is capable of formulating a diffusion model with only four sampling steps, thereby significantly improving the efficiency during inference. The core idea is to corrupt the HQ image towards its LQ counterpart instead of the Gaussian white noise. This strategy can effectively truncate the length of the diffusion model. Extensive experiments on the tasks of image super-resolution and image inpainting have demonstrated the superiority of our proposed method. In addition, more discussion on the limitations of the proposed method can be found in the Appendix. We believe that our work will pave the way for the development of more efficient and effective diffusion models to address the IR problem.